Suppose that {xn} is a convergent sequence and {yn} is such that for any ε>0, there exists M such that |xn−yn|< ϵ for all n≥M. Show that {yn} is a convergent sequence.
1
Expert's answer
2020-06-23T15:44:30-0400

Given that for any "\epsilon>0" , there exists M such that "|x_n-y_n|<\epsilon" for all "n\geq M" ,

"\implies -\epsilon < x_n-y_n < \epsilon \implies x_n-\epsilon<y_n<x_n+\epsilon" for all "n\geq M" .

Now, given "\{x_n\}" is a convergent sequence, "\implies \exist M : |x_n|<\epsilon_1 \ \forall n\geq M_1" .

So, "\exist M_2:|y_n|<max\{\epsilon,\epsilon_1\} \ \forall \ n\geq M_2 = max\{M,M_1\}" .

Hence "\{y_n\}" is a convergent sequence.